An apparatus with a radius of 4.8cm was placed inside a container full of fluorine gas (16.8mol). The temperature of the box is 285.0k and the volume is 142.0 L.

Determine the force pushing down on the apparataus

Force=
Units=
1
Expert's answer
2020-11-18T13:44:12-0500

PV=nRT

P=nRT/V

P= 16.8 × 0.0821×285 / 142

P= 2.768 atm

2.768 atm to Pa = 280467.6Pa

Pressure = Force / Area

Force = Pressure × Area

We assume the apparatus is circular

Force = 280467.6 × π × (0.048)2

= 2030.9 N


Units = Newtons